OpenStudy (anonymous):

The quantity of negative 5 times x minus 10 all divided by negative 5

OpenStudy (theviper):

\[\Large{\frac{-5\times x-10}{5}}\]

OpenStudy (theviper):

\[\Large{\frac{-5x-10}{5}}\]Take 5 as common \[\Large{\frac{5(-x-2)}{5}=-x-2}\]